OpenStudy (anonymous):

4 arithmethic mean between -2 & 23

OpenStudy (anonymous):

What is the question?

OpenStudy (anonymous):

WELCOME TO OPENSTUDY :D

OpenStudy (e.mccormick):

\[\frac{23-2}{2}\]Unless there is more to the question.....

OpenStudy (anonymous):

How about you try using simple examples first?

OpenStudy (anonymous):

let b = 23 and a= -2 then d = b-a/n+1 where n = 4 therefore d = 5 so the arithmetic means are 3,8,13,18

OpenStudy (e.mccormick):

AH HA! Someone that understood the question! It has been to long since I took stats.

OpenStudy (anonymous):

Let the four arithmetic means be A2,A3,A4 and A5. then A1,A2,A3,A4,A5,A6 is our new sequence A1=-2 and A6=23 since An=A1+(n-1)d A6=-2+(6-1)d=23 5d=25 d=5 so An=A(n-1)+d A2=-2+5=3 A3=3+5=8 A4=8+5=13 A5=13=5=18.